OpenStudy (anonymous):

How would I go about solving the following: Lim 2x squared + 5x +2 ---------------- x squared + 2x -8 x-> -2

OpenStudy (shadowfiend):

So it looks a bit better, we're talking about: \[\lim_{x \rightarrow -2} \frac{2x^2 + 5x + 2}{x^2 + 2x - 8}\]

OpenStudy (anonymous):

I started by factoring it and got \[\left[ (x+4)(x+5) \right]\div\left[ (x-4)(x+2) \right]\] The think is that I can not get rid of anything on the bottom. What do I do.